What is the area (sq.ft) of a room if the width is 12ft 3 in and the length is 22ft 3/16 in
1
Expert's answer
2021-10-04T14:02:22-0400

Let's first convert width and length into feet:


"Width=(12\ ft+3\ in\cdot\dfrac{0.083333\ ft}{1\ in})=12.25\ ft,""Length=(22\ ft+\dfrac{3}{16}\ in\cdot\dfrac{0.083333\ ft}{1\ in})=22.016\ ft."

Finally, we can find the area of a room:


"Area=Width\cdot Length,""Area=12.25\ ft\cdot22.016\ ft=269.7\ ft^2."